When 3 dice are thrown, the probability of getting a sum of 3 is
A.
1/6
B.
1/216
C.
1/36
D.
1/120

Found 2 solutions by greenestamps, jim_thompson5910:
Answer by greenestamps(13334)


There are 6 possible outcomes for each die, so the total number of outcomes for three dice (in which order matters) is 6*6*6=216.

There is only one way to get a sum of 3: 1 on each of the three dice.

So the probability of getting a sum of 3 is 1/216.

Answer by jim_thompson5910(35256)


There are 6^3 = 6*6*6 = 36*6 = 216 ways to have three dice land. Only one of those ways is where we have three copies of 1 land face up, so 1+1+1 = 3 as the total sum.

Divide the two values and we get 1/216.

Answer: Choice B) 1/216